Two equal circles are cut out of a rectangular paper of dimensions 16m by 8m. The circles have the maximum diameter possible. what is the approximate area of the paper remaining after the circles have been cut out?
(a) 104
(b) 78
(c) 54
(d) 27
(e) 13

Answer by Cromlix(4381)

Area of paper = 16*8 = 128m^2
Maximum diameter of both circles = 8 metres
Therefore maximum radius for both circles = 4 metres
Area of circle = Pi * radius^2
Area of circle = 50.3 m^2
Two circles = 50.3 * 2 = 100.6 metres^2
Area of paper - Area of two circles.
128m^2 - 100.6m^2
= 27.4 metres^2 of paper left
Answer = (d) 27
Hope this helps.
:-)